Given Lim f(x)=4 and lim g(x)=1/4,



x-›c. x->c



Evaluate the following limits



1.) Lim [-2*f(x)]



x->c



3.) Lim [f(x) +g(x)]



x->c




1
Expert's answer
2022-04-05T02:17:53-0400

"\displaystyle \lim_{\mathclap{x\to c}} {[-2\cdot f(x)]}= -2\cdot\lim_{\mathclap{x\to c}} { f(x)}=-2\cdot4=-8"

"\displaystyle\lim_{\mathclap{x\to c}} {[f(x)+g(x)]} =\displaystyle\lim_{\mathclap{x\to c}} {f(x)}+\lim_{\mathclap{x\to c}} {g(x)} =4+\frac14" "=\frac{17}{4}"